OpenStudy (angelwings996):

Delta H_comb for benzoic acid can be calculated by Delta H=(C_system Delta T)/n, where n is the moles of benzoic acid in the sample and C_system is the heat capacity of the calorimetric system. If the accepted value for the heat of combustion for benzoic acid is 3226 kJ/mol, calculate the heat capacity (C_system) of the calorimetric system

OpenStudy (angelwings996):

Delta T is 3 and moles is 0.0074

OpenStudy (mertsj):

I don't know how to do that. sorry

OpenStudy (angelwings996):

Okay, thanks for looking

OpenStudy (mertsj):

yw. I never heard of that delta H stuff

OpenStudy (anonymous):

\[ \Delta H_{comb}=\frac{C_{system}}{n} \]So you're given \(\Delta H_{comb}\) and they want \(C_{system}\)? Don't you need \(n\)?

OpenStudy (anonymous):

Oops its:\[ \Delta H_{comb}=\frac{C_{system}\Delta T}{n} \]

OpenStudy (angelwings996):

Yes, it says that n is the number of moles and in the before question I had to find it and I got 0.0074 g/mol

OpenStudy (anonymous):

So you need change in temperature?

OpenStudy (angelwings996):

It's just asking to solve for C_system with Delta T being 3, n being the moles : 0.0074, and I think Delta H is 3226 kJ/mol

OpenStudy (anonymous):

If you solve for \(C_{system}\) I think you get:\[ \frac{n\Delta H }{\Delta T}=C_{system} \]

OpenStudy (angelwings996):

Okay, then do I plug in the numbers to get what C is?

OpenStudy (anonymous):

Yes, just gotta look out for units.

OpenStudy (angelwings996):

\[\frac{ 0.0074(3226) }{ 3 } = C _{system}\] like this?

OpenStudy (anonymous):

Yeah,

OpenStudy (angelwings996):

Okay, I got 7.957

OpenStudy (anonymous):

Honestly I am fuzzy on this subject so I can't say if it's correct or not.

OpenStudy (angelwings996):

Okay, well thanks for your help though